# Aligning with Human Judgement: The Role of Pairwise Preference in Large Language Model Evaluators

### v1

@@ Appendix A Additional Related Work
Calibrating LLM Evaluators. Though previous studies have shed light on using LLMs as evaluators, LLMs have been exposed to their sensitivity to prompt designs (Zhao et al., 2021; Zhou et al., 2023b), and even biased from various factors, including but not limited to positional bias (Wang et al., 2023b; Pezeshkpour & Hruschka, 2023), verbosity bias (Saito et al., 2023), and contextual bias (Zhou et al., 2023a). Calibration methods are then proposed to mitigate these biases by criteria optimization (Liu et al., 2023b), context decomposition (Li et al., 2023b), order permutation (Wang et al., 2023b; Zheng et al., 2024a), ensembling (Li et al., 2023a) and batching (Zhou et al., 2023a). However, we show that calibrations are yet insufficient in aligning LLM evaluators with humans in Sec. X.
@@ B.1 Rank Aggregation Likelihood Derivation
$\mathcal{L}(\pi)=\prod_{(i,j)\in C}P(i\succ j|\pi)$
@@
$P(i\succ j|\pi)=\begin{cases}P(i\succ j)&\text{\quad if }\pi(i)<\pi(j)\\ 1-P(i\succ j)&\text{\quad if }\pi(i)>\pi(j)\end{cases}$
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j)*\prod_{i,j:\pi(i)>\pi(j)}(1-P(i\succ j))$
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j)*\prod_{i,j:\pi(i)>\pi(j)}P(j\succ i)$
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j)$
@@ B.2 Transitive Ranking Likelihood Derivation
$=\lambda_{1}P_{\pi_{i},\pi_{i+1}}+\lambda_{2}P_{\pi_{i},\pi_{i+2}}+...+\lambda_{k-1}P_{\pi_{i+k-2},\pi_{i+k-1}}+\lambda_{k-1}P_{\pi_{i+k-1},\pi_{i+k}}$ (6)
@@ B.3 Ranking to Scores Conversion via Quantile Matching
For example, if we assume the score prior distribution is $[10\%,20\%,40\%,20\%,10\%]$ for score 1 to 5, then we can simply convert a ranking into scores by assigning the smallest $10\%$ with score 1, the smallest $10\%-30\%$ with score 2 and so on.
@@ D.1 Evaluation Strategies
Following previous works (Liu et al., 2023a; Zhong et al., 2022; Fu et al., 2023) we used two evaluation strategies, Sample level and Dataset level. For sample-level, the correlation values are calculated across the multiple candidates to the same input and then averaged across all input samples. For dataset-level, the correlation is calculated across the entire dataset.
@@ E.1 Uncertainty threshold for different LLMs
Different base models have different logits distributions. Mistral 7B and Llama2 7B tends to have symmetric and bell-shape distributions, which means large portion of the probabilities lays within the range of $[0.3,0.7]$, as shown in Figure 4. However, as shown in Figure 7, the probability distribution is different for GPT-3.5-turbo. Therefore, their optimal uncertainty thresholds would be different.
Figure 6: Illustration of PairS with a beam size of 2. Heatmap values show $P(A_{i}\succ B_{j})$ by GPT-3.5-turbo. The rest of setting is the same as Figure 4
Figure 7: Empirical evidence for anchor set size selection. The error bars show the mean and standard deviation of the KL divergence between the score distributions of randomly sampled anchor sets (100 repetitions) and the full set. The results suggest that anchor set size of 100 yields relatively small error.
@@ E.2 Theoretical and Empirical Evidence for Choice of Anchor Set Size
The Likert scale score rating can be treated as a five categories classification problem. Therefore the problem of anchor size estimation becomes estimating the sample size for estimating a categorical distribution. This can be calculate by the the method of ‘proportion sample size calculation’ The formula for the sample size calculation for a single proportion is:
$n=\frac{(z^{2}\times p\times(1-p))}{e^{2}}$

### v2

@@ Appendix A Additional Related Work
Calibrating LLM Evaluators. Though previous studies have shed light on using LLMs as evaluators, LLMs have been exposed to their sensitivity to prompt designs (Zhao et al., 2021; Zhou et al., 2023b), and even biased from various factors, including but not limited to positional bias (Wang et al., 2023b; Pezeshkpour & Hruschka, 2023), verbosity bias (Saito et al., 2023), and contextual bias (Zhou et al., 2023a). Calibration methods are then proposed to mitigate these biases by criteria optimization (Liu et al., 2023b), context decomposition (Li et al., 2023b), order permutation (Wang et al., 2023b; Zheng et al., 2024a), ensembling (Li et al., 2023a) and batching (Zhou et al., 2023a). However, we show that calibrations are yet insufficient in aligning LLM evaluators with humans.
@@ B.1 Rank Aggregation Likelihood Derivation
$\mathcal{L}(\pi)=\prod_{(i,j)\in C}P(i\succ j|\pi).$
@@
$P(i\succ j|\pi)=\begin{cases}P(i\succ j)&\text{\quad if }\pi(i)<\pi(j)\\ 1-P(i\succ j)&\text{\quad if }\pi(i)>\pi(j).\end{cases}$
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j)*\prod_{i,j:\pi(i)>\pi(j)}(1-P(i\succ j)).$
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j)*\prod_{i,j:\pi(i)>\pi(j)}P(j\succ i).$
@@
$\mathcal{L}(\pi)=\prod_{i,j:\pi(i)<\pi(j)}P(i\succ j).$
@@ B.2 Transitive Ranking Likelihood Derivation
$=\lambda_{1}P_{\pi_{i},\pi_{i+1}}+\lambda_{2}P_{\pi_{i},\pi_{i+2}}+...+\lambda_{k-1}P_{\pi_{i+k-2},\pi_{i+k-1}}+\lambda_{k-1}P_{\pi_{i+k-1},\pi_{i+k}}.$ (6)
@@ B.3 Ranking to Scores Conversion via Quantile Matching
For example, if we assume the score prior distribution is $[10\%,20\%,40\%,20\%,10\%]$ for scores 1 to 5, then we can simply convert a ranking into scores by assigning the smallest $10\%$ with score 1, the smallest $10\%-30\%$ with score 2 and so on.
@@ D.1 Evaluation Strategies
Following previous works (Liu et al., 2023a; Zhong et al., 2022; Fu et al., 2023) we used two evaluation strategies, Sample level and Dataset level. For the sample level, the correlation values are calculated across the multiple candidates to the same input and then averaged across all input samples. For the dataset level, the correlation is calculated across the entire dataset.
@@ E.1 Uncertainty threshold for different LLMs
Different base models have different logit distributions. Mistral 7B and Llama2 7B tend to have symmetric and bell-shape distributions, which means a large portion of the probabilities lie within the range of $[0.3,0.7]$, as shown in Figure 4. However, as shown in Figure 7, the probability distribution is different for GPT-3.5-turbo. Therefore, their optimal uncertainty thresholds would be different.
Figure 6: Illustration of PairS with a beam size of 2. Heatmap values show $P(A_{i}\succ B_{j})$ by GPT-3.5-turbo. The rest of settings is the same as Figure 4
Figure 7: Empirical evidence for anchor set size selection. The error bars show the mean and standard deviation of the KL divergence between the score distributions of randomly sampled anchor sets (100 repetitions) and the full set. The results suggest that an anchor set size of 100 yields a relatively small error.
@@ E.2 Theoretical and Empirical Evidence for Choice of Anchor Set Size
The Likert scale score rating can be treated as a five-category classification problem. Therefore the problem of anchor size estimation becomes estimating the sample size for estimating a categorical distribution. This can be calculated by the method of ‘proportion sample size calculation’ The formula for the sample size calculation for a single proportion is:
$n=\frac{(z^{2}\times p\times(1-p))}{e^{2}},$